# Float Self-Tagging

### v2

@@ 9. Conclusion
Avoiding heap allocation of floating point numbers has always been a major concern for dynamic and polymorphic languages. This paper presents and evaluates the popular techniques deployed to that end and it presents a new approach, coined self-tagging, that allows attaching type information to 64-bit values at low cost. The core idea is to superimpose tags to bits that appear with high probability in practice. This approach is especially effective to represent IEEE754 floating-point numbers, whose encoding requires all bits of a 64-bit machine word.
Float self-tagging avoids heap-allocation of frequently occurring float values, reducing the strain on the garbage collector and improving performance. Contrary to NaN-boxing and NuN-boxing, which are designed to optimize the handling of floats, self-tagging offers a different trade off that adds no overhead to the handling of other types. Since it does not rely on the specificity of the IEEE754 NaN encoding, self-tagging is highly portable, including to 32-bit architectures.
It is also straightforward to retrofit in implementations that already use object tagging. The only required changes are to reserve one or more tags to float self-tagging and implement the encoding and decoding of the target type. The rest of the runtime stays unchanged.
As shown by the experimental evidence presented in the paper, float self-tagging is a good candidate for languages such as JavaScript, whose specification relies heavily on floats, but has potential applications in any polymorphic language.

### v3

@@ 7. Conclusion
Avoiding heap allocation of floating point numbers has always been a major concern for dynamic and polymorphic languages. This paper presents and evaluates the popular techniques deployed to that end (NaN and NuN-boxing) as well as a new approach, coined self-tagging, that allows representing certain floats without heap allocation.
The core idea is to apply a bitwise transformation that maps the most common floats used in practice to bit patterns that contain the required type information at the correct position. This allows encoding floats in specific ranges as tagged values instead of heap allocated objects, reducing the strain on the garbage collector and improving performance. Contrary to NaN-boxing and NuN-boxing, which are designed to optimize the handling of floats, self-tagging offers a different trade-off that adds no overhead to the handling of other types. Since it does not rely on the specificity of the IEEE754 NaN encoding, self-tagging is highly portable, including to 32-bit architectures.
It is also straightforward to retrofit in implementations that already use object tagging. The only required changes are to reserve one or more tags for float self-tagging and implement an encoding/decoding function for self-tagged floats. The rest of the runtime stays unchanged.
As shown by the experimental evidence presented, float self-tagging is a good candidate for general-purpose languages, such as JavaScript, whose specification relies heavily on floats, but where the performance of other types must not be sacrificed.
@@
8. Data Availability Statement
Data generated and analyzed for this paper is available in a companion artifact (Melançon et al., 2025). This includes results for all microarchitectures listed in Section 4, some of which were omitted in this paper for lack of space. The artifact also contains all source code used to run the experiments.